# Capturing a Moving Target by Two Robots in the F2F Model

### v1

@@ Abstract.
In our analysis, we are mindful of the cost of changing direction of movement, and show how to accomplish the capture of the target with at most three direction changes (turns). Key words and phrases. Autonomous robot, Capture time, Competitive ratio, F2F (Face-to-Face), Knowledge model, Oblivious target, Searcher, Speed, Turn.
@@ 1.1. Preliminaries and Notation
We consider algorithms under various constraints reflecting the knowledge the two robots have about the initial distance of the target from the origin, its direction of movement (away or toward), and the speed of the moving target. We will study algorithms under different knowledge models. The FullKnowledge (FK) model refers to two mobile agents that have complete knowledge of the initial distance, direction of movement, and the speed of the moving target. The NoSpeed model refers to no knowledge of the speed of the target. The NoDistance (ND) model refers to no knowledge of the initial distance of the target. The NoKnowledge (NK) model refers to the case when the robots have no knowledge of the speed and the initial distance of the target.
@@ 1.2. Related Work
Our work is strongly influenced by and builds upon the results of (coleman2022line, ), where the authors analyzed the competitive ratio under four knowledge models: FullKnowledge, NoSpeed, NoDistance, and NoKnowledge. While their analysis focuses on a single searcher, we extend their work to two searchers operating in the F2F communication model. Additionally, the tightness of the main lower bound for the NoSpeed model with a single searcher was recently established in (coleman2024lineiwoca, ).
@@ 1.3.3. No Speed Model
$d$ Algorithm 8 Away $1+\frac{16\log^{2}\left(\frac{1}{1-v}\right)}{(1-v)^{8}}$
$d$ Algorithm 9 Toward $3$
@@ 1.3.4. No Knowledge Model
None Algorithm 10 Away $O\left(M^{12}(\log^{2}M)/d\right)$
@@
Note that the number of turns (changes of direction) in all algorithms is at most 3, with the exception of Algorithms 4 and 6. All proofs missing from the main text can be found in the appendix.
@@ 2.2. Target moving toward the origin
We prove the following result.
@@ 3.1. Target moving away from the origin
The competitive ratio of Algorithm 4 is at most $\frac{(v+3)^{2}}{(1-v)^{2}}$, where $a=\frac{2(1+v)}{1-v}$.
@@
$R_{1}$ moves in one direction with speed $u$.
$R_{2}$ moves in the other direction with speed $u$.
@@
(Theorem 3.3) For robot $R_{1}$ to reach the target, it needs time
$\frac{d}{u-v}.$
At this point, $R_{2}$ will be at a distance $\frac{2du}{u-v}$ from the target. In order for $R_{1}$ to catch $R_{2}$, it needs time
@@
At this point, both robots will be at a distance of $\frac{2du+2duv}{(u-v)(1-u)}$ from the exit. Thus, the time needed for both robots to reach the target is
@@
The optimal speed is $u=\frac{3v+1}{3+v}$, and the competitive ratio is:
@@ 3.2. Target Moving Toward the Origin
1:The mobile target starts at an unknown distance $d$ and escapes in a direction unknown to the searchers, moving with speed $v$;
2:The two searchers start at the origin of the real line and agree on a speed $u<1$;
3:The two searchers search in opposite directions with speed $u$;
4:When one of the searchers meets the mobile target, it reverses direction and changes its speed to $1$;
5:When this searcher catches up to the other searcher, they both reverse direction and go together to meet the moving target;
@@
Clearly, this value makes sense only when $v<\frac{1}{3}$ because when $u>0$, it is only in this range. If we plug in $u=\frac{1-3v}{3-v}$ into the formula for $CR(u,v)$ above, we obtain
@@
Given that $10\leq u<$ 1, we note that the value $u=\frac{1-3v}{3-v}$ does not make sense when $v\geq\frac{1}{3}$. In this case, we can employ the waiting algorithm in which both robots wait at the origin for the moving target to arrive. The competitive ratio of the waiting algorithm is $1+\frac{1}{v}$.
@@ 4.1. Target Moving Away from the Origin
We prove the following result.
@@
$\frac{8}{(1-v)^{6}}+32\frac{\log^{2}\left(\frac{1}{1-v}\right)}{(1-v)^{7}}+\frac{128\cdot\log^{2}\left(\frac{1}{1-v}\right)}{(1-v)^{8}}.$
@@
$\leq 2^{\sum_{j=0}^{i+1}f_{j}}\cdot d\cdot 4^{i+1}.$
If the target is captured at iteration $i$ by one of the robots, the competitive ratio $CR$ is given by:
$CR$ $=\frac{\sum_{j=0}^{i-1}x_{j}+\frac{d+v\sum_{j=0}^{i-1}x_{j}}{u_{i}-v}+\frac{2\sum_{j=0}^{i-1}u_{j}x_{j}+2u_{i}\left(\frac{d+v\sum_{j=0}^{i-1}x_{j}}{u_{i}-v}\right)}{1-u_{i}}}{\frac{d}{1-v}}$
$+\frac{\frac{2\sum_{j=0}^{i-1}u_{j}x_{j}+2u_{i}\left(\frac{d+v\sum_{j=0}^{i-1}x_{j}}{u_{i}-v}\right)+2v\sum_{j=0}^{i-1}u_{j}x_{j}+2vu_{i}\left(\frac{d+v\sum_{j=0}^{i-1}x_{j}}{u_{i}-v}\right)}{(1-u_{i})(1-v)}}{\frac{d}{1-v}}$
@@
$+\frac{u_{i}^{2}v\sum_{j=0}^{i-1}x_{j}+du_{i}v+2u_{i}\sum_{j=0}^{i-1}u_{j}x_{j}-2v\sum_{j=0}^{i-1}u_{j}x_{j}+2u_{i}d}{d(1-u_{i})(u_{i}-v)}$
$+\frac{2vu_{i}\sum_{j=0}^{i-1}x_{j}+2u_{i}\sum_{j=0}^{i-1}u_{j}x_{j}-2v\sum_{j=0}^{i-1}u_{j}x_{j}+2u_{i}d+2u_{i}v\sum_{j=0}^{i-1}x_{j}}{d(1-u_{i})(u_{i}-v)}.$
After simplifying the above expressions, we reach the final competitive ratio:
$CR\leq\frac{8}{(1-v)^{6}}+32\frac{\log^{2}\left(\frac{1}{1-v}\right)}{(1-v)^{7}}+\frac{128\cdot\log^{2}\left(\frac{1}{1-v}\right)}{(1-v)^{8}}.$
@@ 5. No Knowledge Model
In this section, we consider the No Knowledge model and distinguish the cases where the target is moving away or toward the origin. {EK: Should we mention that when they zigzag separately the competitive ratio (CR) is $\frac{1}{(1-v)^{8}}$?}

### v2

@@ Abstract.
In our analysis, we are mindful of the cost of changing direction of movement, and show how to accomplish the capture of the target with at most three direction changes (turns).
Key words and phrases. Autonomous robot, Capture time, Competitive ratio, F2F (Face-to-Face), Knowledge model, Oblivious target, Searcher, Speed, Turn.
@@ 1.1. Preliminaries and Notation
We consider algorithms under various constraints reflecting the knowledge the two robots have about the initial distance of the target from the origin, its direction of movement (away or toward), and the speed of the moving target. We will study algorithms under different knowledge models. The FullKnowledge (FK) model refers to two mobile agents that have complete knowledge of the initial distance, direction of movement, and the speed of the moving target. The NoSpeed model refers to no knowledge of the speed of the target. The NoDistance (ND) model refers to no knowledge of the initial distance of the target. The NoKnowledge (NK) model refers to the case when the robots have no knowledge of the speed or the initial distance of the target.
@@ 1.2. Related Work
Our work is strongly influenced by and builds upon the results of (coleman2022line, ), where the authors analyzed the competitive ratio under four knowledge models: FullKnowledge, NoSpeed, NoDistance, and NoKnowledgeOfSpeedAndDistance. While their analysis focuses on a single searcher, we extend their work to two searchers operating in the F2F communication model. Additionally, the tightness of the main lower bound for the NoSpeed model with a single searcher was recently established in (coleman2024lineiwoca, ).
@@ 1.3.2. No Distance Model
$v$ Algorithm 5 Away $\frac{(v+3)^{2}}{(1-v)^{2}}$
$v$ Algorithm 6 Toward $(\frac{v-3}{v+1})^{2}$ if $v\geq 3$
@@ 1.3.3. No Speed Model
$d$ Algorithm 8 Away $O\left(M^{8}(\log^{2}M)/d\right)$
$d$ Algorithm 9 Toward 33
@@ 1.3.4. No Knowledge Model
None Algorithm 10 Away $O\left(M^{10}(\log^{2}M)/d\right)$
@@
Note that the number of turns (changes of direction) in all algorithms is at most 3, with the exception of Algorithms 4 and 6.
@@ 2.1. Target moving away from the origin
Proof.
(Theorem 2.1) Robots $R_{1}$ and $R_{2}$ move in the same direction for a distance of $\frac{d}{1-v}$. If neither of them catches the target during this time, then the target will be ahead of the robots by a distance of $\frac{2d}{1-v}$. To cover this remaining distance, they need an additional time of $\frac{\frac{2d}{1-v}}{1-v}$.
Therefore, the total time required for both robots to capture the target is:
$\frac{d}{1-v}+\frac{\frac{2d}{1-v}}{1-v}=\frac{3d-dv}{(1-v)^{2}}.$
This yields the competitive ratio:
$CR=\frac{\frac{3d-dv}{(1-v)^{2}}}{\frac{d}{1-v}}=\frac{3-v}{1-v}.$
This proves Theorem 2.1. ∎
@@
Proof.
(Theorem 2.2) If both robots $R_{1}$ and $R_{2}$ are at the origin and the moving target is at a distance $d$ away from the origin, then any of the two robots needs $\frac{d}{1-v}$ to catch up with the target. Assume $R_{1}$ reaches point $\frac{d}{1-v}$, then the adversary would place the target at $\frac{d}{1-v}$. At this point, $R_{2}$ catches the target and $R_{1}$ would be far from the target by $\frac{2d}{1-v}$. Thus, for $R_{1}$ to reach the target, it needs time $\frac{\frac{2d}{1-v}}{1-v}=\frac{2d}{(1-v)^{2}}$. The competitive ratio is then:
$CR=\frac{\frac{2d}{(1-v)^{2}}+\frac{d}{1-v}}{\frac{d}{1-v}}=\frac{3-v}{1-v}.$
This proves Theorem 2.2. ∎
@@ 2.2. Target moving toward the origin
Proof.
$\frac{d}{v+1}+\frac{2d}{(v+1)^{2}}=\frac{3d+dv}{(v+1)^{2}}.$
The competitive ratio is then:
$CR=\frac{\frac{3d+dv}{(v+1)^{2}}}{\frac{d}{v+1}}=\frac{3+v}{1+v}.$
This proves Theorem 2.3. ∎
@@
We prove the following results.
@@
Proof.
$CR=\frac{\frac{d}{v}}{\frac{d}{v+1}}=\frac{v+1}{v}.$
This proves Theorem 2.4. ∎
@@
Proof.
  • Case 1: If $v\geq 1$, then clearly $CR\geq 1+\frac{1}{v}$.
This proves Theorem 2.5. ∎
@@ 3.1. Target moving away from the origin
Setting $a=\frac{2(1+v)}{1-v}$, the competitive ratio of Algorithm 4 is at most $\frac{(v+3)^{2}}{(1-v)^{2}}$.
Proof.
(Theorem 3.1) If the target is at distance $d$ from the origin, then eventually either $R_{1}$ or $R_{2}$ will capture the target. Without loss of generality, assume that $R_{1}$ captures the target. The worst-case scenario occurs if, during iteration k1k-1, $R_{1}$ just misses the target. Specifically, when $R_{1}$ reaches $x_{k-1}$, the target is at distance $x_{k-1}+\epsilon$. Thus, for $R_{1}$ to capture the target, it requires time:
(3) $\frac{x_{k}+vx_{k}}{1-v}.$
Using equations (1), (2), and (3), the competitive ratio is given by:
$=\frac{2x_{k}+d+2\sum_{i=0}^{k-1}x_{i}}{d}$
(4) $=\frac{2\sum_{i=0}^{k-1}x_{i}+2x_{k}}{d}+1.$
To analyze further, assume that in round $k$, the robot captures the target. Then, in round k1k-1, the following must hold:
$\frac{d+\left(2a^{0}+\dots+2a^{k-2}\right)v}{1-v}\geq a^{k-1},$
$\implies d>a^{k-1}-va^{k-1}-2v\frac{a^{k-1}-1}{a-1},$
$\implies d>\frac{a^{k}-a^{k-1}-va^{k}+va^{k-1}-2va^{k-1}+2v}{a-1},$
(5) $\implies d>\frac{a^{k}-a^{k-1}-va^{k}-va^{k-1}+2v}{a-1}.$
Thus, the competitive ratio becomes:
(6) $CR$ $=\frac{2ax_{k}-2}{a^{k}-a^{k-1}-va^{k}-va^{k-1}+2v}+1\$$=\frac{2a^{2}}{a-1-av-v}+1.$
$f^{\prime}(a)$ $=\frac{4a(a-1-av-v)-2a^{2}(1-v)}{(a-1-av-v)^{2}}$
$=\frac{4a^{2}-4a-4va^{2}-4av-2a^{2}(1-v)}{(a-1-av-v)^{2}}$
(7) $=\frac{2a^{2}-2va^{2}-4a-4av}{(a-1-av-v)^{2}}.$
$CR$ $=1+\frac{8(1+v)^{2}}{(1-v)^{2}\cdot(a-1-av-v)}$
$=\frac{8(1+v)^{2}}{(1-v)(1-v^{2})}$
$=\frac{8(1+v)}{(1-v)^{2}}+1$
(8) $=\frac{(v+3)^{2}}{(1-v)^{2}}.$
This completes the proof of Theorem 3.1. ∎
@@
Proof.
(Theorem 3.2) Let us assume that the target was detected by one of the robots at iteration $k$. Then at the end of iteration k1k-1, both robots would have completed 2(k1)2(k-1) turns. At iteration $k$, one of the robots would make a turn to meet the other robot at the origin, since the other robot already caught the target. Thus, the total number of turns would be 2(k1)+1=2k12(k-1)+1=2k-1.
In order to find the number of turns in terms of $d$ and $v$, we assume that the target is caught by one of the robots at iteration $k$. Then, at iteration k1k-1, and based on Equation 5, we have:
$\frac{a^{k-1}(a-1-av-v)}{a}\leq d,$
which implies the following:
$a^{k-1}$ $\leq\frac{ad}{a-1-av-v},$
$\leq\frac{\frac{2d(1+v)}{1-v}}{\frac{2+2v-1+v-2v-2v^{2}-v+v^{2}}{1-v}},$
$=\frac{2d(1+v)}{1-v^{2}},$
$=\frac{2d}{1-v}.$
Taking the logarithm, we obtain:
$k\leq 1+\log\left(\frac{2d}{1-v}\right).$
Thus, the total number of turns would be at most:
$2k-1=1+2\log\left(\frac{2d}{1-v}\right).$
This proves Theorem 3.2. ∎
@@
[…]
@@
(Theorem 3.3) According to Algorithm 5, for robot $R_{1}$ to reach the target, it needs time
$\frac{d}{u-v},$
where $u=\frac{3v+1}{3+v}$. At this point, $R_{2}$ will be at a distance $\frac{2du}{u-v}$ from the target. In order for $R_{1}$ to catch $R_{2}$, it needs time
@@
At this point, both robots will be at a distance of $\frac{2du+2duv}{(u-v)(1-u)}$ from the target. Thus, the time needed for both robots to reach the target is
@@
The optimal speed is $u=\frac{3v+1}{3+v}$, and the competitive ratio can be shown to be:
@@ 3.2. Target Moving Toward the Origin
Proof.
(Proof of Theorem 3.5)
If the target is at distance $d$ away from the origin, then eventually either $R_{1}$ or $R_{2}$ will capture the target. If we assume that $R_{1}$ captures the target, then the worst-case scenario occurs when, during iteration k1k-1, $R_{1}$ just misses the target. In other words, when $R_{1}$ reaches $x_{k-1}$, the target would be at distance $x_{k-1}+\epsilon$. Thus, in order to capture the target, $R_{1}$ needs time:
(11) $\frac{x_{k}-vx_{k}}{1+v}$
Thus, using equations 9, 10, and $1$, the competitive ratio is as follows:
$\leq\frac{2\sum_{i=0}^{k-1}x_{i}+2x_{k}+d}{d}$
$=1+\frac{2x_{k+1}-2}{d(a-1)}$
If we assume that in round $k$, the robot captures the target, then in round k1k-1 we should have the following:
$\frac{d-\left(2a^{0}+\dots+2a^{k-2}\right)v}{1+v}\geq a^{k-1}$
$\implies d>a^{k-1}+va^{k-1}+2v\frac{a^{k-1}-1}{a-1}$
Thus, the competitive ratio is as follows:
$CR$ $\leq 1+\frac{2a^{2}\left(\frac{2v+ad-d}{a+av+v-1}\right)-2}{d(a-1)}$
$\leq 1+\frac{4va^{2}+2da^{3}-2da^{2}-2a-2av-2v+2}{d(a-1)(a+av+v-1)}$
$\leq 1+\frac{4av+2da^{2}-2+2v}{d(a+av+v-1)}$
$\leq 1+\lim_{d\to\infty}\frac{4av+2da^{2}-2+2v}{d(a+av+v-1)}$
$=1+\frac{2a^{2}}{a+av+v-1}$
If we set $f(a)=\frac{2a^{2}}{a+av+v-1}$, then
$f^{\prime}(a)$ $=\frac{4a(a+av+v-1)-2a^{2}(1+v)}{(a+av+v-1)^{2}}$
$=\frac{4a^{2}+4va^{2}+4av-4a-2a^{2}-2va^{2}}{(a+av+v-1)^{2}}$
$\leq\frac{2a^{2}+2va^{2}+4av-4a}{(a+av+v-1)^{2}}$
Setting $f^{\prime}(a)=0$ gives 2a+2av+4v4=02a+2av+4v-4=0. Thus, $a=\frac{2(1-v)}{1+v}$.
Thus, the competitive ratio would be:
$CR$ $\leq 1+\frac{\frac{8-16v+8v^{2}}{(1+v)^{2}}}{\frac{2-2v^{2}+v+v^{2}-1-v}{1+v}}$
$\leq 1+\frac{8v^{2}-6v+8}{(1+v)(1-v^{2})}$
$=1+\frac{8(1-v)}{(1+v)^{2}}$
This proves Theorem 3.5. ∎
@@
[…]
3:if $R_{1}$ reaches the target then
4:  It switches direction until it catches $R_{2}$.
5:  Both robots switch direction to meet the target.
6:else
7:  if $R_{2}$ reaches the target then
8:   It switches direction until it catches $R_{1}$.
9:   Both robots switch direction to meet the target.
@@
If we plug in $u=\frac{1-3v}{3-v}$ into the formula for $CR(u,v)$ above, we obtain
@@
Given that $0\leq u<1$, we note that the value $u=\frac{1-3v}{3-v}$ does not make sense when $v\geq\frac{1}{3}$. In this case, we can employ the waiting algorithm in which both robots wait at the origin for the moving target to arrive. The competitive ratio of the waiting algorithm is $1+\frac{1}{v}$.
@@ 4.1. Target Moving Away from the Origin
We prove the following results.
@@
$\frac{5}{2(1-v)^{6}}+\frac{22\cdot\log^{2}\left(\frac{1}{1-v}\right)}{(1-v)^{8}}.$
@@
$\leq 2^{\sum_{j=0}^{i+1}f_{j}}\cdot 4^{i+1}.$
If the target is captured at iteration $i$ by one of the robots, then the competitive ratio $CR$ becomes as follows:
$CR$ $=\frac{\sum_{j=0}^{i-1}x_{j}+\frac{d+v\sum_{j=0}^{i-1}x_{j}}{u_{i}-v}+\frac{2u_{i}\sum_{j=0}^{i-1}x_{j}+2u_{i}\left(\frac{d+v\sum_{j=0}^{i-1}x_{j}}{u_{i}-v}\right)}{1-u_{i}}}{\frac{d}{1-v}}$
$+\frac{\frac{2u_{i}\sum_{j=0}^{i-1}x_{j}+2u_{i}\left(\frac{d+v\sum_{j=0}^{i-1}x_{j}}{u_{i}-v}\right)+2u_{i}v\sum_{j=0}^{i-1}x_{j}+2vu_{i}\left(\frac{d+v\sum_{j=0}^{i-1}x_{j}}{u_{i}-v}\right)}{(1-u_{i})(1-v)}}{\frac{d}{1-v}}$
@@
$+\frac{u_{i}^{2}v\sum_{j=0}^{i-1}x_{j}+du_{i}v+4u_{i}^{2}\sum_{j=0}^{i-1}x_{j}+4u_{i}d}{d(1-u_{i})(u_{i}-v)}.$
Since the target is detected at iteration $i$, we have $2^{2^{i-1}}<\frac{1}{1-v}$, and since $2^{i-1}\leq\log\left(\frac{1}{1-v}\right)$, we get
$4^{i+1}\leq 16\log^{2}\left(\frac{1}{1-v}\right).$
Considering $a_{i}=1+\frac{1}{2^{2^{i}}}$, we have the following:
$u_{i}-v$ $>u_{i}-v_{i}=\frac{v_{i}}{2^{2^{i}}}\geq\frac{(1-v)^{2}}{2}.$
And
$1-u_{i}$ $=1-\frac{v_{i}}{2^{2^{i}}}-v_{i}$
$=1-v_{i}\left(1+\frac{1}{2^{2^{i}}}\right)$
$=1-\left(1-2^{-2^{i}}\right)\left(1+\frac{1}{2^{2^{i}}}\right)$
$=1-1-\frac{1}{2^{2^{i}}}+2^{-2^{i}}+\frac{1}{2^{2^{i+1}}}$
$=\frac{1}{2^{2^{i+1}}}\geq(1-v)^{4}.$
Next, we have
$\sum_{j=0}^{i-1}x_{j}$ $\leq\frac{x_{i}(u_{i}-v_{i})}{v_{i}}$
$\leq 2(u_{i}-v_{i})2^{\sum_{j=0}^{i}f_{j}}\cdot 4^{i}$
$\leq 2\frac{v_{i}}{2^{2^{i}}}\cdot 2^{2^{i+1}}\cdot 4^{i}$
$\leq 8\cdot\left(\frac{1}{1-v}\right)^{2}\cdot\log^{2}\left(\frac{1}{1-v}\right).$
The competitive ratio becomes:
$CR\leq\frac{5}{2(1-v)^{6}}+\frac{22\cdot\log^{2}\left(\frac{1}{1-v}\right)}{(1-v)^{8}}.$
@@ 4.2. Target Moving Toward the Origin
Proof of Theorem 4.3.
In the worst-case scenario, if the target’s speed is very small, the robots may not catch up to the target immediately. In this case, the robots will move a distance $d$ in one direction. At this point, they will be away from the target by a distance of 2ddv2d-dv, where $v$ is the speed of the target.
The competitive ratio (CR) can then be calculated as follows:
[…]
$=\frac{d+\frac{2d-dv}{1+v}}{\frac{d}{1+v}}$
$=3.$
Hence, the competitive ratio is $3$, which proves Theorem 4.3. ∎
@@
Proof of Theorem 4.4.
To prove the lower bound, both points $d$ and d-d must be visited by the robots. If only one of these two points, say d-d, is visited while the other point $d$ is not, the adversary could place the target at $d$. In this case, the robots would fail to catch the target efficiently.
Similarly, if the target is placed at d-d and one of the robots visits $d$ first, then by the time the robot reaches $d$, the target would have moved at least $dv$. Following this argument, the competitive ratio would be calculated as follows:
[…]
$=\frac{d+\frac{2d-dv}{1+v}}{\frac{d}{1+v}}$
$=3.$
@@ 5. No Knowledge Model
In this section, we consider the No Knowledge model and distinguish the cases where the target is moving away or toward the origin.